# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### v6

@@ A.2 Financial Institutions in the Cities
Notes: This table summarizes the number of depositors and the amount of deposits in the savings banks by occupation. These statistics were measured on 31st December 1919. Panel A and B list the statistics for Tokyo City and Kyobashi Word, respectively. Source: Tokyo City Office (1922a, pp. 864–865).
@@
Notes: This map shows the location of the pawnshops and moneylenders in Tsukishima. The bullseye indicates the pawnshop (shichiya). The black circle shows the money lenders (kanekashi gyō). The white circle shows the rickshaw station (kuruma yado). Source: Department of Health, Ministry of the Interior 1923c, tenth map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
Figure A.2: Pawnshops and Money Lenders
@@ A.3 Development of Tsukishima
Notes: Tsukishima contains Ishikawajima, Tsukudajima, Shin-tsukudajima, and Tsukishima Lots 1–3. This map does not illustrate the northernmost part of Ishikawa jima (see Figure 3 for this part). “Ferries” shows the Tsukishima to Kyobashi Ward sea routes: Tsukuda, Tsukishima, and Kachidoki routes (right to left). “Aioi bridge” links Shin tsukudajima and Fukagawa ward. Source: Department of Health, Ministry of the Interior 1923c, 15th map. The author modified and adjusted the tone using Adobe Photoshop 24.7.0.
@@ A.4 Housings in Tsukishima
Note: This table summarizes the housing types in Tsukishima surveyed on 30th June 1919. Source: Department of Health, Ministry of the Interior 1923a, pp. 69–76.
Figure A.4: Photographs of the Tenements in Tsukishima
@@
Note: Figures A.4a and A.4b show the two- and one-story tenements in Tsukishima, respectively. Source: Department of Health, Ministry of the Interior 1923c, figures 62–63; 66.
@@ A.5 Machinery Factories in Tsukishima
Table A.5: Industrial Structure and Social Classes in Tsukishima: Evidence from Tsukishima Primary School Survey
@@
Notes: This table summarizes the statistics from the complete survey on the 4th-6th graders in Tsukishima1st and 2nd Primary Schools in July 1919. Panel A summarizes the heads’ industries, and Panel B sorts the “social classes” of the heads in the manufacturing sector, respectively. The report does not specify the definition of each “social class”. Source: Department of Health, Ministry of the Interior 1923a, pp. 386–387.
@@
Notes: Panel A summarizes the number of machinery factories by the scale and type of the factory. Statistics are from the manufacturing census conducted in the Tsukishima Survey (1st November 1920). The threshold of the scale (i.e., 15 workers) is based on the application of the Factory Act enacted in 1916. Panel B sorted the heads by their skill level and job type. Statistics are obtained from the Complete Survey on the 4th-6th graders in Tsukishima1st and 2nd Primary Schools in July 1919. Source: Panel A: Department of Health, Ministry of the Interior 1923a, pp. 386–387; 410–414. Panel B: Department of Health, Ministry of the Interior 1923a, p. 149.
@@ B.1 Tsukishima Survey: The First Social Survey in Japan
First, I summarize the investigators, research questions, and subjects of the Tsukishima Survey. The person responsible for the Tsukishima Survey was Iwasaburō Takano (hereafter “Takano”), a professor of social statistics and social policy at the Law School of Tokyo Imperial University. Takano became a professor of economics at Tokyo Imperial University in 1919, and became the director of the OISR in 1920. Takano placed great importance on gathering objective statistics about workers through social surveys (Sekiya 1970, p. 33). He had the experience of conducting the first household survey in Japan in 1916.808080This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo” and was inspired by Charles Booth (Miyoshi 1980, p. 33). Although the original documents of this survey were preserved at Tokyo Imperial University, those were destroyed in the Great Kantō Earthquake (Takano 1933, pp. 727–728). More precisely, before this household survey, Takano conducted a field survey in a poor area of Tokyo when he was a student at Tokyo Imperial University. The findings were reported in April 1894 as an exercise entitled “East London in Tokyo” (Kawai 1981b, p. 3). Takano has been a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.818181The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No. 172. Initially, 34 members were appointed, and the first meeting was held on July 8. The reason for the establishment of this committee was based on the need for a policy response regarding the establishment and institutionalization of various investigation organizations and the penetration of the capitalist economy during the Meiji period. Another reason is the requirement for a public health response to the rising mortality rate, exemplified by the tuberculosis epidemic (Kawai 1981a, pp. 10–13). In 1917, the Seventh Section evolved into the Section on Urban and Rural Sanitary Conditions (Kawai 1981a, pp. 12–13). On October 22, 1918, Takano submitted a proposal to the subcommittee for a field survey of urban sanitary conditions, which was approved. The survey aimed to determine the housing conditions, household budget status, and health status (child health, production, stillbirths, deaths, and diseases) in the areas where skilled-worker families congregated (Sekiya 1970, p. 6). Initially, the survey area was planned to be Yanagibashi Yokogawa-chō in Honjo Ward. However, at a committee meeting on November 23, the survey site was changed to Tsukishima in Kyobashi Ward, which had more skilled worker households than in Yokogawa-chō (Kawai 1980, pp. 55–59). [^82] Takeda (2015, pp. 75–76) provides more details on how the survey site was changed. Takano was looking for a commissioned researcher to conduct the survey and asked Sakuzō Yoshino, a colleague at Tokyo Imperial University, to recommend a suitable person. Yoshino approached his former student Kotora Tanahashi, who had become a member of the fraternity club, a mutual aid organization for workers, to accept the investigator position. The diary written by Kotora Tanahashi (“Tanahashi Kotora Diary”) is in the possession of the OISR, and Takeda used the diary as the basis for her analysis of the circumstances surrounding the change of survey location. Tanahashi eventually recommended Yoshitsuru Yamana, who was his classmate in high school and Tokyo Imperial University, as a researcher and considered Tsukishima to be the best research area because of the ease of selecting survey households using the fraternity’s network (Takeda 2015, pp. 76–77). On October 26, Tanahashi and Yamana, together with the Kyobashi Branch of the fraternity club, conducted an on-site inspection of Tsukishima. Based on this inspection, Tanahashi and Yamana approached Takano about locating the survey station in Tsukishima. Finally, Takano and Yamana conducted an on-site inspection of Tsukishima on October 30 and decided on Tsukishima as the survey site.
@@ 10. Elementary school survey July 1919 2
Notes: The years/months listed in the first column (“Periods of the field studies”) are based on the description in the Department of Health, Ministry of the Interior (1923a, pp. 10–17). Figures listed in the second column (“Chapter in the published report” ) are based on the table of contents in the Department of Health, Ministry of the Interior (1923a, pp. 1–12; 1923b, pp.1–2). “A2” indicates the Appendix 2. Sources: Department of Health, Ministry of the Interior (1923a, 1923b, 1923c).
@@
Notes: This photograph shows an example of the official report of the Tsukishima Survey published by the Department of Health, Ministry of the Interior, in 1923. Source: Department of Health, Ministry of the Interior 1923b, title page. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
Figure B.2: Official Published Report of the Tsukishima Survey
Finally, I summarize the format of the published report and the unpublished materials. The results of the Tsukishima Survey were officially summarized in one report published in December 1921 by the Sanitary Bureau of the Ministry of the Interior and in two appendices (Figure B.2 is an example of the front cover). The report was published under the name “Report on Field Survey in Tsukishima, Kyobashi, Tokyo,” and consists of four sections. The first section is a comprehensive review by Takano and outlines the survey. The second section summarizes the living conditions of the workers and was written by Gonda (partly by Yamana). Hoshino wrote the third section and summarized the sanitary conditions of workers. In the fourth section, Yamana summarizes the labor situation in Tsukishima. As for the two appendices, one contains statistical tables related to the above report, and the other organizes social maps and photographs. The correspondence between the survey items and the report is summarized in Table B.1. The Ministry of Home Affairs has already published these materials, which are available at the Gonda Yasunosuke Library of the OISR and the University of Tokyo Library.848484 In the OISR, these are included in the Archives of Gonda Yasunosuke (7-2; 7-3; 7-4). In addition, a second manuscript of the Tsukishima Survey report exists. This is the manuscript of the second series, which was preserved in the home of Gonda Hayao, the second son of Gonda Yasunosuke, and was discovered by Terade Koji. The second series includes a summary of the results of a household survey of elementary school teachers’ households in Tokyo, conducted as a supplementary survey to the Tsukishima Survey (see Terade (1982) for details). The manuscript of this second series was later donated to the OISR and is now available as part of the Gonda Yasunosuke Collection at the Institute (The OISR, Archives of Gonda Yasunosuke (8-12; 8-13; 8-14; 8-15)).
@@ B.2 Original Micro-spreadsheets from the Household Budget Survey in Tsukishima Survey
Notes: This photograph depicts an example page from the household expenses account of the THBS. The right page documents the details of costs. The left page includes the details of revenue and receiving and making payments on the goods (actual things, including gifts). Source: The THBS, account #6, 15th February 1919. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
@@ B.3 Trimming
Note: Figure B.4aB.4b, and B.4c show the observations in the daily, semi-monthly, and adjusted monthly panel datasets, respectively. For example, a marker in Figure B.4a indicates an observation in a specific year-month-date cell. Strictly, semi-monthly, and monthly datasets include both panel and cross-sectional units because aggregation shall trim fractions in the edges of the survey period. Figure B.4b shows 35-panel units and 5 cross-sectional units (#29; #31-33; #35). Figure B.4c also shows 35-panel units and 5 cross-sectional units (#29; #31-33;#35). Source: Created by the author using the THBS sample.
@@
Figure B.4 summarizes the structure of the 40 THBS households. Figure B.4aB.4b, and B.4c shows the observations in the daily, semi-monthly, and adjusted monthly datasets, respectively. As explained, there are five cross-sectional units (#29; #31-33; #35) in the semi-monthly and adjusted monthly datasets. Therefore, there are 35 units with a panel structure. After excluding a unit without complete income information (#37), I get the 34 THBS households as the analytical sample for my panel data analysis.
Regarding the time dimension, I trim all the observations before 12th January 1919 and after 11th November 1919 because the number of surveyed households is substantially small. For example, Figure B.4b shows that the number of cross-sectional units is three or less in both periods, causing substantial fluctuations in the net income. Consequently, my analytical sample includes 34 households measured between 12th January and 11th November 1919. The average number of semi-months per unit is $8.8$ (Std. Dev. $=5.0$). The characteristics of the 34 THBS households are summarized in detail in Section 3.
@@ B.4 Testing the Potential Influence of the Unbalancedness
The results from Probit models are reported. Robust standard errors are in brackets. Notes: The dependent variable is an indicator variable that takes one if the household has shorter time-series observations. Column 1 uses the median of the number of days per unit as the threshold. The result from the regression using the average number of days per unit ($144$ days) as the threshold is identical to that listed in column 1. Column 2 uses the 75 percentile of the number of days per unit as the threshold. Column 3 uses the 25 percentile of the number of days per unit as the threshold. All family size variables are the time-series average in each unit over the sample period. The proportion of children aged 0–5 years (%) is used as the reference group. Wald $\chi^{2}$ statistics $p$-values for the null of the zero-slope hypothesis are reported in the sixth row. The results from the expanded regressions including the head’s age and earnings are materially similar. The null of the zero-slope hypothesis is not rejected in all the specifications at the conventional level (not reported).
@@ B.5 Comparing Distributions of the Household Size
Figure B.5 illustrates the household size distributions measured in the census and THBS sample. Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$. This supports the evidence that the THBS sample has a similar household size distribution to the population of Tsukishima (Section 3.1). As described in Section 3.2, while the THBS sample does not cover households with one person and with a very large number of people, it covers most of the classes in the household sizes with reasonable approximation to the population distribution. [^87] The number of households with one people is $362$, accounting for only $6$% of the total number of households. Similarly, the number of households with nine or more people is $294$, accounting for only $5$% (Section 3.2).
@@
Notes: This figure shows the share of the household in each household size category to the total number of households in Tsukishima measured in the 1920 Population Census and the THBS samples. Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$. The share in each household size measured in the census is used as the theoretical probability in calculating $\chi^{2}$ test statistic (Pearson 1900). Sources: Created by the author using the Tokyo City Office (1922a, pp. 262–283) and the THBS dataset.
Figure B.5: Distribution of the Household Size:
Population Census v. the THBS sample

### v7

@@ A.2 Financial Institutions in the Cities
Notes:
1. Panel A: The statistics on postal savings were measured at the end of March 1921. The statistics on savings banks were measured at the end of 1920.
@@
Sources: Panel A: Tokyo City Office (1923, pp. 969; 977); Osaka City Office (1922, pp. 6(61); 6(67)); Kyoto Prefecture (1922, p. 257); Kyoto City Office (1926, pp. 188); Nagoya City Office (1922, pp. 286; 290). Panel B: Tokyo City Office (1923, pp. 888–889); Osaka City Office (1922, pp. 6(68)–(69)); Kyoto City Office (1922, p. 191); Tokyo City Statistics Division (1926b, p. 12).
@@
Notes: This table summarizes the number of depositors and the amount of deposits in the savings banks by occupation. In the document, the miscellaneous category is not clearly defined but named as “Miscellaneous (zatsu gyō)”. These statistics were measured on 31st December 1919. Panel A and B list the statistics for Tokyo City and Kyobashi Word, respectively.
Source: Tokyo City Office (1922a, pp. 864–865).
@@
Notes: This map shows the location of the pawnshops and moneylenders in Tsukishima. The bullseye indicates the pawnshop (shichiya). The black circle shows the money lenders (kanekashi gyō). The white circle shows the rickshaw station (kuruma yado).
Source: Department of Health, Ministry of the Interior 1923c, tenth map. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
@@ A.3 Development of Tsukishima
Notes: Tsukishima contains Ishikawajima, Tsukudajima, Shin-tsukudajima, and Tsukishima Lots 1–3. This map does not illustrate the northernmost part of Ishikawa jima (see Figure 3 for this part). “Ferries” shows the Tsukishima to Kyobashi Ward sea routes: Tsukuda, Tsukishima, and Kachidoki routes (right to left). “Aioi bridge” links Shin tsukudajima and Fukagawa ward.
Source: Department of Health, Ministry of the Interior 1923c, 15th map. The author modified and adjusted the tone using Adobe Photoshop 24.7.0.
@@ A.4 Housings in Tsukishima
Note: This table summarizes the housing types in Tsukishima surveyed on 30th June 1919.
Source: Department of Health, Ministry of the Interior 1923a, pp. 69–76.
@@
Note: Figures A.4a and A.4b show the two- and one-story tenements in Tsukishima, respectively.
Source: Department of Health, Ministry of the Interior 1923c, figures 62–63; 66.
@@ A.5 Machinery Factories in Tsukishima
Table A.5: Industrial Structure and Social Classes in Tsukishima:
Evidence from Tsukishima Primary School Survey
@@
Notes: This table summarizes the statistics from the complete survey on the 4th-6th graders in Tsukishima1st and 2nd Primary Schools in July 1919. Panel A summarizes the heads’ industries, and Panel B sorts the “social classes” of the heads in the manufacturing sector, respectively. The report does not specify the definition of each “social class”.
Source: Department of Health, Ministry of the Interior 1923a, pp. 386–387.
@@
Notes: Panel A summarizes the number of machinery factories by the scale and type of the factory. Statistics are from the manufacturing census conducted in the Tsukishima Survey (1st November 1920). The threshold of the scale (i.e., 15 workers) is based on the application of the Factory Act enacted in 1916.
Panel B sorted the heads by their skill level and job type. Statistics are obtained from the Complete Survey on the 4th-6th graders in Tsukishima1st and 2nd Primary Schools in July 1919.
Source: Panel A: Department of Health, Ministry of the Interior 1923a, pp. 386–387; 410–414. Panel B: Department of Health, Ministry of the Interior 1923a, p. 149.
@@ B.1 Tsukishima Survey: The First Social Survey in Japan
First, I summarize the investigators, research questions, and subjects of the Tsukishima Survey. The person responsible for the Tsukishima Survey was Iwasaburō Takano (hereafter “Takano”), a professor of social statistics and social policy at the Law School of Tokyo Imperial University. Takano became a professor of economics at Tokyo Imperial University in 1919, and became the director of the OISR in 1920. Takano placed great importance on gathering objective statistics about workers through social surveys (Sekiya 1970, p. 33). He had the experience of conducting the first household survey in Japan in 1916.939393This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo” and was inspired by Charles Booth (Miyoshi 1980, p. 33). Although the original documents of this survey were preserved at Tokyo Imperial University, those were destroyed in the Great Kantō Earthquake (Takano 1933, pp. 727–728). More precisely, before this household survey, Takano conducted a field survey in a poor area of Tokyo when he was a student at Tokyo Imperial University. The findings were reported in April 1894 as an exercise entitled “East London in Tokyo” (Kawai 1981b, p. 3). Takano has been a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.949494The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No. 172. Initially, 34 members were appointed, and the first meeting was held on July 8. The reason for the establishment of this committee was based on the need for a policy response regarding the establishment and institutionalization of various investigation organizations and the penetration of the capitalist economy during the Meiji period. Another reason is the requirement for a public health response to the rising mortality rate, exemplified by the tuberculosis epidemic (Kawai 1981a, pp. 10–13). In 1917, the Seventh Section evolved into the Section on Urban and Rural Sanitary Conditions (Kawai 1981a, pp. 12–13). On October 22, 1918, Takano submitted a proposal to the subcommittee for a field survey of urban sanitary conditions, which was approved. The survey aimed to determine the housing conditions, household budget status, and health status (child health, production, stillbirths, deaths, and diseases) in the areas where skilled-worker families congregated (Sekiya 1970, p. 6). Initially, the survey area was planned to be Yanagibashi Yokogawa-chō in Honjo Ward. However, at a committee meeting on November 23, the survey site was changed to Tsukishima in Kyobashi Ward, which had more skilled worker households than in Yokogawa-chō (Kawai 1980, pp. 55–59). [^95] Takeda (2015, pp. 75–76) provides more details on how the survey site was changed. Takano was looking for a commissioned researcher to conduct the survey and asked Sakuzō Yoshino, a colleague at Tokyo Imperial University, to recommend a suitable person. Yoshino approached his former student Kotora Tanahashi, who had become a member of the fraternity club, a mutual aid organization for workers, to accept the investigator position. The diary written by Kotora Tanahashi (“Tanahashi Kotora Diary”) is in the possession of the OISR, and Takeda used the diary as the basis for her analysis of the circumstances surrounding the change of survey location. Tanahashi eventually recommended Yoshitsuru Yamana, who was his classmate in high school and Tokyo Imperial University, as a researcher and considered Tsukishima to be the best research area because of the ease of selecting survey households using the fraternity’s network (Takeda 2015, pp. 76–77). On October 26, Tanahashi and Yamana, together with the Kyobashi Branch of the fraternity club, conducted an on-site inspection of Tsukishima. Based on this inspection, Tanahashi and Yamana approached Takano about locating the survey station in Tsukishima. Finally, Takano and Yamana conducted an on-site inspection of Tsukishima on October 30 and decided on Tsukishima as the survey site.
@@ 10. Elementary school survey July 1919 2
Notes: The years/months listed in the first column (“Periods of the field studies”) are based on the description in the Department of Health, Ministry of the Interior (1923a, pp. 10–17). Figures listed in the second column (“Chapter in the published report” ) are based on the table of contents in the Department of Health, Ministry of the Interior (1923a, pp. 1–12; 1923b, pp.1–2). “A2” indicates the Appendix 2.
Sources: Department of Health, Ministry of the Interior (1923a, 1923b, 1923c).
@@
Notes: This photograph shows an example of the official report of the Tsukishima Survey published by the Department of Health, Ministry of the Interior, in 1923.
Source: Department of Health, Ministry of the Interior 1923b, title page. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
Finally, I summarize the format of the published report and the unpublished materials. The results of the Tsukishima Survey were officially summarized in one report published in December 1921 by the Sanitary Bureau of the Ministry of the Interior and in two appendices (Figure B.2 is an example of the front cover). The report was published under the name “Report on Field Survey in Tsukishima, Kyobashi, Tokyo,” and consists of four sections. The first section is a comprehensive review by Takano and outlines the survey. The second section summarizes the living conditions of the workers and was written by Gonda (partly by Yamana). Hoshino wrote the third section and summarized the sanitary conditions of workers. In the fourth section, Yamana summarizes the labor situation in Tsukishima. As for the two appendices, one contains statistical tables related to the above report, and the other organizes social maps and photographs. The correspondence between the survey items and the report is summarized in Table B.1. The Ministry of Home Affairs has already published these materials, which are available at the Gonda Yasunosuke Library of the OISR and the University of Tokyo Library.979797 In the OISR, these are included in the Archives of Gonda Yasunosuke (7-2; 7-3; 7-4). In addition, a second manuscript of the Tsukishima Survey report exists. This is the manuscript of the second series, which was preserved in the home of Gonda Hayao, the second son of Gonda Yasunosuke, and was discovered by Terade Koji. The second series includes a summary of the results of a household survey of elementary school teachers’ households in Tokyo, conducted as a supplementary survey to the Tsukishima Survey (see Terade (1982) for details). The manuscript of this second series was later donated to the OISR and is now available as part of the Gonda Yasunosuke Collection at the Institute (The OISR, Archives of Gonda Yasunosuke (8-12; 8-13; 8-14; 8-15)).
@@ B.2 Original Micro-spreadsheets from the Household Budget Survey in Tsukishima Survey
Notes: This photograph depicts an example page from the household expenses account of the THBS. The right page documents the details of costs. The left page includes the details of revenue and receiving and making payments on the goods (actual things, including gifts).
Source: The THBS, account #6, 15th February 1919. The tone was adjusted by the author using Adobe Photoshop 24.7.0.
@@ B.3 Trimming
Note: Figure B.4aB.4b, and B.4c show the observations in the daily, semi-monthly, and adjusted monthly panel datasets, respectively. For example, a marker in Figure B.4a indicates an observation in a specific year-month-date cell. Strictly, semi-monthly, and monthly datasets include both panel and cross-sectional units because aggregation shall trim fractions in the edges of the survey period. Figure B.4b shows 35-panel units and 5 cross-sectional units (#29; #31-33; #35). Figure B.4c shows 28-panel units and 6 cross-sectional units (#20; #25; #26; #28; #34; #40). 5 cross-sectional units in the semi-monthly panel do not appear in this figure (#29; #31-33; #35). Note that unit #30 having two semi-monthly cells is not included because those cells are separated into different adjusted months (February and March, respectively).
Source: Created by the author using the THBS sample.
@@
Figure B.4 summarizes the structure of the 40 THBS households. Figures B.4a and B.4b show the observations in the daily and semi-monthly datasets, respectively. As explained, there are five cross-sectional units (#29; #31-33; #35) in the semi-monthly dataset. Therefore, there are 35 units with a panel structure. After excluding two units without complete information on income and credit purchases (#37; #39), I get the 33 THBS households as the analytical sample for my panel data analysis.
Regarding the time dimension, I trimmed all the observations before 12th January 1919 and after 11th November 1919 because the number of surveyed households is substantially small. For example, Figure B.4b shows that the number of cross-sectional units is three or less in both periods, causing substantial fluctuations in the net income. Consequently, my analytical sample includes 33 households measured between 12th January and 11th November 1919. The average number of semi-months per unit is $8.8$ (Std. Dev. $=5.0$). The characteristics of the 33 THBS households are summarized in detail in Section 3.
Regarding the adjusted monthly series, we further exclude several units that have insufficient observations for constructing the monthly panel: 26 THBS households are included in my adjusted monthly panel dataset. Figure B.4c shows the observations in the adjusted monthly dataset.
@@ B.4 Testing the Potential Influence of the Unbalancedness
The results from Probit models are reported. Robust standard errors are in brackets.
Notes: The dependent variable is an indicator variable that takes one if the household has shorter time-series observations. Column 1 uses the median of the number of days per unit as the threshold. The result from the regression using the average number of days per unit ($144$ days) as the threshold is identical to that listed in column 1. Column 2 uses the 75 percentile of the number of days per unit as the threshold. Column 3 uses the 25 percentile of the number of days per unit as the threshold. All family size variables are the time-series average in each unit over the sample period. The proportion of children aged 0–5 years (%) is used as the reference group. Wald $\chi^{2}$ statistics $p$-values for the null of the zero-slope hypothesis are reported in the sixth row. The results from the expanded regressions including the head’s age and earnings are materially similar. The null of the zero-slope hypothesis is not rejected in all the specifications at the conventional level (not reported).
@@ B.5 Comparing Distributions of the Household Size
Figure B.5 illustrates the household size distributions measured in the census and THBS sample. Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$. This supports the evidence that the THBS sample has a similar household size distribution to the population of Tsukishima (Section 3.1). As described in Section 3.2, while the THBS sample does not cover households with one person and with a very large number of people, it covers most of the classes in the household sizes with reasonable approximation to the population distribution.100100100The number of households with one people is $362$, accounting for only $6$% of the total number of households. Similarly, the number of households with nine or more people is $294$, accounting for only $5$% (Section 3.2).
@@
Notes: This figure shows the share of the household in each household size category to the total number of households in Tsukishima measured in the 1920 Population Census and the THBS samples. Pearson chi-squared test does not reject the null hypothesis of the equality of distributions with $p$-value of $0.948$. The share in each household size measured in the census is used as the theoretical probability in calculating $\chi^{2}$ test statistic (Pearson 1900).
Sources: Created by the author using the Tokyo City Office (1922a, pp. 262–283) and the THBS dataset.